OpenStudy (anonymous):

Choose the value for x that makes the open sentence true. 8 + 3 • x = 22 + x2

OpenStudy (blurbendy):

is that 8 + 3 = 22 + x^2

OpenStudy (anonymous):

oops its should be 2 squared plus x squared

OpenStudy (blurbendy):

8 + 3x = 2^2 + x^2 ?

OpenStudy (blurbendy):

We can turn that into a quadratic equation: x^2 -3x + (8 + 4) = x^2 - 3x + 12 I think you'll have to use the quadratic formula here.

OpenStudy (anonymous):

whats the value of x

OpenStudy (blurbendy):

You need to use the quadratic formula to figure it out

OpenStudy (anonymous):

i dont know how

OpenStudy (blurbendy):

 We had x^2 - 3x + 12, so a = 1 b = -3 c = 12 Plug them into the formula to solve for x